Verify that the point P(acosθ,bsinθ) lies on the ellipse
x^2/ a^2+ y^2/ b^2= 1,
where a and b are the semi-major and semi-minor axes respectively of the ellipse . Find the gradient of the tangent to the curve at P and show that the equation of the normal at P is axsinθ−by cosθ = (a^2 −b^2)sinθ cosθ. If P is not on the axes and if the normal at P passes through the point B(0,b), Show that a^2 > 2b^2. If further, the tangent at P meets the y-axis at Q, show that
|BQ| =
a^2/ b^2
.
1
Expert's answer
2020-07-01T18:39:32-0400

equation of ellipse x2/a2 + y2/b2 = 1

we have points P(acos"\theta" + bsin"\theta" )

put the points in the equation of the ellipse

take L.H.S

(acos"\theta")2/a2 + (bsin"\theta")2/b2

a2cos2"\theta" /a2 + b2sin2 "\theta"/b2

cos2 "\theta" + sin2"\theta"

=1

hence the points are lies in the ellipse


the equation of normal is at p is ax sin"\theta" - by cos"\theta" = (a2-b2)sin"\theta" cos"\theta"

the gradient of the equtation is dy/dx

so the gradient is

dy/dx( ax sin"\theta") - dy/dx(by cos"\theta" )= dy/dx {(a2-b2)sin"\theta"cos"\theta"}

asin"\theta" -0=0 is the gradient of the tangent to the curve at P


equation of the normal is

axsin"\theta" -bycos"\theta" = (a2-b2)sin"\theta" cos"\theta" at point B(0,b)

a(0)sin"\theta" - b(b)cos"\theta" = (a2-b2)sin"\theta" cos"\theta"

0-b2cos"\theta" =(a2-b2)sin"\theta" cos"\theta"

hence a2>2b2


if the tangent P meets the y axis at Q

than |BQ| = a2/b2

the point B (0,b) similarly at y axis the point Q (a,0)

so the the point BQ(a,b)

hence |BQ| = a2/b2